OpenStudy (anonymous):

4x+y=-10 in function form

OpenStudy (anonymous):

y=-4x-10

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

4x + y = -10 y = -10 - 4x y = f(x) f(x) = -4x -10